can you made fraction into decimal?
1
Expert's answer
2011-12-09T09:07:34-0500
Fractions and decimals represent the same things: numbers that are not whole numbers. Decimal is a fraction which denomanator is power of 10< we just do not write it. For example, 0.5=5/10. To change a fraction into a decimal, divide the numerator (top number) by the denominator (bottom number). Example: What is 5/8 as a decimal?get your calculator and type in "5 / 8 ="
The answer should be 0.625. If you want to do it without calculator you can do it as following. Consider example.
Let's write 9/25 as a decimal.
25 is not a power of 10, but we can easily make it a power of 10 -- we can make it 100 -- by multiplying it by 4.& We must also, then, multiply the numerator by 4. Thus we have& 9/25=9*4/25*4=36/100=0.36
One more& example. Express 1/3 as a decimal. There is no way to multiply 3 to become 10 or 100 or any "1 followed by 0s", but we can calculate an approximate decimal by choosing to multiply by, say, 333. Multiply top and bottom by 333:
1/3=333/999. Now, 999 is nearly 1,000, so let us write down 333 with the decimal point 3 spaces from the right (because 1,000 has 3 zeros): 0.333 (accurate to only 3 decimal places)

In conclusion, to convert a fraction into decimal you should make it's denominator a power of ten by multiplying or dividing numerator and denominator by certain number.